OpenStudy (anonymous):

Help Please. 4^2x+3=1

OpenStudy (anonymous):

I need help badly in couple of questions, someone please help!!

jimthompson5910 (jim_thompson5910):

is the problem \[\Large 4^{2x}+3=1\] OR is it \[\Large 4^{2x+3}=1\]

OpenStudy (anonymous):

yeah that second one

jimthompson5910 (jim_thompson5910):

Hint: \[\Large 4^{2x+3}=1\] \[\Large 4^{2x+3}=4^0\]

OpenStudy (anonymous):

so we are not using that 1?

jimthompson5910 (jim_thompson5910):

we are, just a different form of it 4^0 is the same as 1

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

the bases are the same (4) so the exponents must be the same

jimthompson5910 (jim_thompson5910):

so 2x + 3 = 0

OpenStudy (anonymous):

yup i Got the answer thanks :)

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

can you please help me with couple of them too?

jimthompson5910 (jim_thompson5910):

sure a few more

OpenStudy (anonymous):

ok here is another one

OpenStudy (anonymous):

\[10^{-3x}\times10^{x}=\frac{ 1 }{ 10 }\]

jimthompson5910 (jim_thompson5910):

\[10^{-3x}\times10^{x}=\frac{ 1 }{ 10 }\] \[10^{-3x}\times10^{x}=10^{-1}\] \[10^{-3x+x}=10^{-1}\] I'll let you finish

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

I need help with couple more pleaseeee!!

OpenStudy (anonymous):

\[64\times16^{-3x}=16^{3x-2}\]

jimthompson5910 (jim_thompson5910):

-3x + x = -1 -2x = -1 x = 1/2 good

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

\[64\times16^{-3x}=16^{3x-2}\] \[(4^3)\times(4^2)^{-3x}=(4^2)^{3x-2}\] \[(4^3)\times4^{2(-3x)}=(4)^{2(3x-2)}\] \[(4^3)\times4^{-6x}=(4)^{6x-4}\] \[4^{3+(-6x)}=(4)^{6x-4}\] \[4^{3-6x}=(4)^{6x-4}\]

OpenStudy (anonymous):

I got an answer: \[x=\frac{ 7 }{ 12 }\]

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

\[\frac{ 81^{3n+2} }{ 243^{-n} }=3^{4}\]

OpenStudy (anonymous):

oh maybe I can do this one let me try first :)

jimthompson5910 (jim_thompson5910):

alright, tell me what you get

OpenStudy (anonymous):

I got \[-\frac{ 4 }{ 17 }\]

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[81\times9^{-2b-2}=27\]

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh I got that one \[-\frac{ 3 }{ 4 }\]

jimthompson5910 (jim_thompson5910):

good, that's correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 6\end{matrix}\right)^{3x+2}\times216^{3x}=\frac{ 1 }{ 216 }\]

jimthompson5910 (jim_thompson5910):

and what did you get

OpenStudy (anonymous):

I m having trouble starting it but still trying

jimthompson5910 (jim_thompson5910):

\[\Large \left(\frac{1}{6}\right)^{3x+2}\times216^{3x}=\frac{ 1 }{ 216 }\] \[\Large \left(6^{-1}\right)^{3x+2}\times(6^3)^{3x}=6^{-3}\] \[\Large \left(6\right)^{-(3x+2)}\times(6)^{9x}=6^{-3}\] \[\Large 6^{-(3x+2)+9x}=6^{-3}\]

OpenStudy (anonymous):

oh I only messed up with one sign